OpenStudy (anonymous):

Box A has a volume of 48 cubic meters. Box B is similar to box A. To create Box B, Box A's dimensions were doubled. What is the volume of Box B? A.24 m3 B.96 m3 C.288 m3 D.384 m3

OpenStudy (anonymous):

@rishavraj

rishavraj (rishavraj):

volume = l*b*h volume of A = 48 when dimensions were doubled then volume = 2l*2b*2h

rishavraj (rishavraj):

@mmrachel done???

OpenStudy (anonymous):

what do i have to do?

OpenStudy (anonymous):

@rishavraj

rishavraj (rishavraj):

as i told l*b*h = 48 so wht would be 2l*2b*2h?????

OpenStudy (anonymous):

i dont know the length or the base or the heighth

rishavraj (rishavraj):

see like l*b*h = 48 so 2l*2b*2h = 8(l*b*h) and u know the value of l**b*h

OpenStudy (anonymous):

so it would be 8x8x8

rishavraj (rishavraj):

-_- 8*48

OpenStudy (anonymous):

oh sorry so it would be 384?

rishavraj (rishavraj):

yup :))))

OpenStudy (anonymous):

yay! thanks

OpenStudy (anonymous):

again!

rishavraj (rishavraj):

yw again lol :P